OpenStudy (barterrr):

I need help ASAP!!! ( i don't really remember geometry) Question 1.) What is the relationship between the sine and cosine of complementary angles? How do we know it is true?

OpenStudy (rayep12):

@sammixboo

OpenStudy (mertsj):

Co-functions of complementary angles are equal.

OpenStudy (rayep12):

oh hey Mertsj

OpenStudy (barterrr):

@mertsj And do you mind explaining how it is true please!

OpenStudy (mertsj):



OpenStudy (mertsj):

A and B are complementary angles. sine and cosine are co-functions \[\sin A=\frac{BC}{AB}\] \[\cos B=\frac{BC}{AB}\]